Fermi National Laboratory


Accelerator Update

Wednesday June 30
The day shift began with operators, vacuum technicians, and Tevatron (TeV) experts investigating a possible TeV ground fault, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:39 AM, a Booster RF station (BRF9) tripped off on a LCW conductivity problem. The station reset okay.

At 9:45 AM, Pbar stopped stacking so that the Main Injector (MI) could conduct a slipstacking study.

At 10:13 AM, the Pbar Accumulator suffered from a vacuum burst at A150.

Operators began a TeV checkout at 10:56 AM. TeV experts then began a study period.

MI experts stopped their slipstacking studies at 12:04 PM.

Operations put the TeV through a dry squeeze at 12:53 PM.

At 3:15 PM, a Collider Detector at Fermilab (CDF) experimenter called to report that they had lost power to their computers.

At 3:27 PM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum burst. Operations reset a vacuum pump.

TeV experts completed their studies at 4:34 PM and Operations put the TeV into shot setup.

At 7:16 PM, Operations reported that they had loaded the protons and antiprotons in the TeV when the TeV beam aborted due to a sextupole power supply (T:SF) tripping off. Operators investigated and discovered there was a spike on T:SF just before the trip. Two power supply experts were called in.

Operations prepared the D-Zero (D0) collision hall for access at 8:42 PM.

Thursday July 1
The midnight shift began with TeV experts working on the T:SF power supply, with Pbar stacking, and with MiniBooNE taking beam.

TeV experts reported at 12:32 AM that they had completed their work on T:SF.

Operations put the TeV through a dry squeeze at 12:37 AM.

Operations loaded the TeV with a 12x0 study-store at 2:50 AM.

At 3:46 AM, a Booster power supply (ORBMP) tripped off due to bad vacuum. Operators investigated and discovered a vacuum pump with a failed power supply. Operators replaced the power supply at 4:26 AM. Operations resumed sending beam to MiniBooNE.

Operations put the TeV into shot setup at 4:43 AM.

Operations established store 3610 at 7:33 AM with an initial luminosity of 36.33E30.

Operations established beam to the Meson MTest experiment at 10:11 AM.

Operations halted Pbar stacking at 10:39 AM so that techs could access the Pre-Vault area to work on shunt equipment for a quadrupole (QS928).

A Linac expert began conducting wire-scans at 10:44 AM. This halted MiniBooNE beam for about forty-five minutes.

At 11:34 AM, the Booster ORBMP power supply tripped off again due to bad vacuum. Both ORBMP and the vacuum pumps reset okay. Techs changed some control modules.

Booster beam resumed at 12 PM.

At 2:06 PM, Operations reported that a Linac RF station (KRF5) had tripped off a half dozen times in five minutes. Operators reset the station each time.

At 3 PM, Operations reported that a Booster RF station (BRF12) had tripped off four times in the last fifteen minutes. It had tripped off several times earlier in the shift. An expert investigated and reported that the Power Amplifier (PA) for this station was failing. He made adjustments to help it last a little longer.

BRF12 tripped off at 4:26 PM. It reset okay.

The MTest experiment finished with beam for the night at 9:27 PM.

The Meson MCenter experiment requested beam at 11:47 PM.

At 11:50 PM, the Booster lost seven RF stations (BRF1-8) due to the east anode power supply tripping off.

Friday July 2
The midnight shift began with Operations monitoring store 3610, with Pbar stacking, and with the MiniBooNE and MCenter experiments taking beam.

Operations terminated store 3610 at 2:01 AM and then put the TeV into shot setup at 2:49 AM.

KRF5 suffered a number of trips between 4:03 AM and 4:18 AM.

At 4:43 AM, operators had to put the TeV shot setup on hold after the third antiproton injection. The TeV sequencer was acting as if this was a Mixed Mode shot (antiprotons coming from both Pbar and the Recycler). Operations contacted an expert and then reset the sequencer.

Operations established store 3612 at 5:50 AM with an initial luminosity of 62.46E30.

CDF experimenters asked Operations to resolve a colliding beam problem at 6:23 AM. Operators were able to resolve this problem and CDF was able to turn their detector on at 6:57 AM.

At 7:43 AM, a Linac RF station (KRF7) tripped off and would not reset do to a failed power supply. Operators could not find a spare. A Linac expert began changing out the supply at 8:16 AM.

Plans for Today and this Weekend
The plans for today and this weekend are to stack and store. The Recycler will stash antiprotons after Operations establishes the next store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/2/2004   email Fermilab